# The Orientation Boundary for Step-Duplicating Recursors: Mechanized Impossibility, Escape, and Certification

### 19 Mar 2026 v7

@@ 4.2 Affine / linear constructor-local measures
Machine-checked in Meta/StepDuplicatingSchema.lean.
Proof sketch. Fix $b=n=\mathrm{base}$ and pump $s$. Write
@@
Thus the generic affine contradiction bound is exact for this family. Machine-checked as affineThresholdMeasure_exact_cutoff.
@@
with $\beta_{w},\gamma_{w}\geq 1$. If $M$ has unbounded attainable values, then $M$ does not strictly orient the duplicating step uniformly. The same conclusion follows from either a positive successor pump or a positive wrap/base pump, exactly as in Theorem 4.2. Machine-checked in Meta/QuadraticBarrier.lean as no_quadratic_counter_orients_dup_step_of_unbounded together with the successor- and wrap-pump corollaries.
Proof sketch. Fix $b=n=\mathrm{base}$ and pump the step argument $s$. The rewrite source is
@@
formalized in Meta/QuadraticCrossTermBarrier.lean as CrossTermBoundedAtBase. Then $M$ does not strictly orient the duplicating step uniformly. In particular, the same conclusion follows from either a positive successor pump or a positive wrap/base pump together with this bounded-coupling condition. Machine-checked in Meta/QuadraticCrossTermBarrier.lean as no_cross_quadratic_orients_dup_step_of_unbounded and the corresponding KO7 specializations.
@@
Proof sketch. Fix $b=n=\mathrm{base}$ and pump the step argument $s$. The source coefficient in the pump variable becomes $\gamma_{r}+\zeta_{r}\,M(\mathrm{succ}(\mathrm{base}))$, while the target coefficient becomes $\beta_{w}+\gamma_{w}(\gamma_{r}+\zeta_{r}\,M(\mathrm{base}))$. The bounded-coupling hypothesis says the target coefficient still dominates the source coefficient by at least one. Once the pumped step argument exceeds the frozen constant part, the target overtakes the source and strict decrease is impossible. This pushes the barrier one step beyond the pure counter-square class: some step-counter coupling is still excluded, provided the wrapper gain remains dominant at the base point. $\square$
@@
where $\mathcal{T}$ is a finite table of multilinear monomials, each variable appearing at most once per monomial. Assume $M$ has unbounded attainable values and that, after freezing $b=\mathrm{base}$ and comparing the source counter $\mathrm{succ}(\mathrm{base})$ with the target counter $\mathrm{base}$, the total coefficient of the pumped step variable on the source side is still dominated by the wrapper-side coefficient. Then $M$ does not strictly orient the duplicating step uniformly. The same conclusion follows from either a positive successor pump or a positive wrap/base pump under the same dominance condition. Machine-checked in Meta/MultilinearBarrier.lean as no_multilinear_orients_dup_step_of_unbounded together with the successor-pump, wrap-pump, and KO7 corollaries.
Proof sketch. Freeze $b=\mathrm{base}$ and pump the step argument $s$. Each multilinear monomial becomes affine in $M(s)$ once the other arguments are frozen, so the entire finite monomial table collapses to a constant part plus a frozen coefficient times $M(s)$ on both the source and target sides. The explicit dominance condition says the wrapper-side coefficient still exceeds the frozen source coefficient by at least one. Once $M(s)$ exceeds the frozen constant part, the target overtakes the source and strict decrease becomes impossible. This sharpens the nonlinear boundary: a large finite multilinear family still fails as long as its pumped-variable contribution remains frozen and dominated in this sense. $\square$
@@
Let $(T,\mathrm{base},\mathrm{succ},\mathrm{wrap},\mathrm{recur})$ be a step-duplicating schema. Let $M:T\to\mathbb{N}$ satisfy constructor-local affine equations for $\mathrm{succ}$ and $\mathrm{wrap}$, and let the recursor be given by a finite polynomial table in the three tracked scalar arguments $M(b)$, $M(s)$, and $M(n)$, with arbitrary natural exponents allowed in each monomial. Assume $M$ has unbounded attainable values and that, after freezing $b=\mathrm{base}$ and comparing the source counter $\mathrm{succ}(\mathrm{base})$ with the target counter $\mathrm{base}$, the frozen target-side polynomial eventually dominates the frozen source-side polynomial as the pumped step value grows. Then $M$ does not strictly orient the duplicating step uniformly. The same conclusion follows from either a certified successor pump or a certified wrap pump under the same frozen dominance condition. Equivalently, any successful orienter inside this bounded direct-polynomial family must violate the frozen base-dominance condition. Machine-checked in Meta/PolynomialBarrierGeneral.lean as no_polynomial_orients_dup_step_of_unbounded, no_global_step_orientation_polynomial_of_succ_pump, and polynomial_escape_requires_failure_of_base_dominance.
@@
Proof sketch. Freeze the non-pumped arguments at the base point and compare the source counter $\mathrm{succ}(\mathrm{base})$ with the target counter $\mathrm{base}$. Each monomial then becomes a one-variable polynomial in the pumped step value, possibly with repeated occurrences of that variable. The explicit dominance hypothesis says that beyond some threshold the frozen target-side polynomial overtakes the frozen source-side polynomial. Pump the step value beyond that threshold. The target is then at least the source, so strict decrease fails. This extends the freeze-and-pump argument beyond the multilinear regime: repeated variables are allowed, but only if the resulting one-variable growth remains dominated at the frozen base point. $\square$
@@
If $M$ has unbounded attainable values, then $M$ does not strictly orient the duplicating step uniformly. The same conclusion follows from either a positive successor drift or a positive left-wrapper drift. Machine-checked in Meta/MaxBarrier.lean as no_max_orients_dup_step_of_unbounded together with the successor-pump, wrap-pump, and KO7 corollaries.
Proof sketch. Pump the step argument until the step branch $\gamma_{r}+M(s)$ dominates both the base branch and the counter branch inside the recursor. At that point the source becomes $\alpha_{r}+\gamma_{r}+M(s)$, while the target wraps the same visible branch again through a max-plus wrapper. Because the wrapper keeps a visible copy of the pumped branch, the target is at least the source plus one of the wrapper contributions, so strict decrease fails. This adds a different aggregation regime to the barrier stack: the obstruction persists even when aggregation is by $\max$ rather than by addition. $\square$
@@
Any direct arctic-style interpretation with a distinguished finite primary projection satisfying the max-plus interface of Theorem 4.7 is blocked by the same argument. In particular, the KO7 root relation admits no such arctic primary-projection orienter under the corresponding unbounded or certified pump hypotheses. Machine-checked in Meta/ArcticBarrier.lean as no_arctic_primary_orients_dup_step_of_unbounded and the KO7 corollary no_global_step_orientation_arctic_primary_of_unbounded.
Proof sketch. Forget the arctic carrier and read only the distinguished finite primary coordinate. By hypothesis that projection already satisfies the max-plus interface, so Theorem 4.7 applies to it directly. The arctic corollary is therefore a tool-facing instantiation of the scalar max barrier, not a separate max-algebra proof. $\square$
@@
Fix any finite dimension $d\geq 1$. Let $M:T\to\mathbb{N}^{d}$ be a vector-valued constructor-local measure on a step-duplicating schema, ordered by strict componentwise decrease. Assume one tracked coordinate of $M$ satisfies the affine equations of Theorem 4.2 with positive wrapper coefficients and an unbounded pump in that tracked coordinate. Then the duplicating step cannot be strictly oriented componentwise by $M$. In particular, the same conclusion follows if the tracked coordinate admits either a positive successor pump or a positive wrap/base pump. Machine-checked in Meta/MatrixBarrierD.lean as no_matrixD_orients_dup_step_of_componentwise_pump and the corresponding KO7 specialization. The earlier dimension-$2$ development is retained in Meta/MatrixBarrier2.lean as the concrete low-dimensional worked instance.
Proof sketch. Strict componentwise orientation implies strict decrease in every coordinate, hence in the tracked one. Project the vector-valued measure to that tracked coordinate and apply Theorem 4.2. This yields a contradiction immediately. The result remains intentionally narrow: it covers arbitrary fixed dimension under strict componentwise order, but still only through one tracked affine coordinate, not through arbitrary mixed matrix interactions. $\square$
@@
Let $M:T\to\mathbb{N}^{2}$ be a pair-valued constructor-local measure on a step-duplicating schema, ordered lexicographically by the first component. Assume the first component of $M$ satisfies the affine equations of Theorem 4.2 with positive wrapper coefficients and an unbounded pump in that first component. Then the duplicating step cannot be strictly oriented lexicographically by $M$. In particular, the same conclusion follows if the first component admits either a positive successor pump or a positive wrap/base pump. Machine-checked in Meta/MatrixBarrierLex.lean as no_matrix2_lex_orients_dup_step_of_unbounded_primary and the corresponding KO7 specializations.
Proof sketch. Lexicographic decrease implies that the first component of the target is at most the first component of the source. Pump the tracked first component one unit beyond the scalar affine bound. The first component of the duplicating-step target then exceeds the first component of the source, so lexicographic decrease is impossible regardless of what the second component does. This remains a tracked-primary result, not a theorem about arbitrary mixed matrix interpretations. $\square$
@@
Let $M:T\to\mathbb{N}^{2}$ be a pair-valued constructor-local measure on a step-duplicating schema, ordered by strict componentwise decrease. Assume each constructor acts by a full mixed $2\times 2$ linear map on the two coordinates, and that the two column sums of each such map agree. Then the aggregate coordinate sum $M_{1}+M_{2}$ is an affine scalar measure. If that aggregate sum has an unbounded pump and the aggregate wrapper coefficients are positive, then the duplicating step cannot be strictly oriented componentwise by $M$. Machine-checked in Meta/MatrixBarrierMix2.lean as no_matrixMix2_orients_dup_step_of_sum_pump and the corresponding KO7 specialization.
Proof sketch. Balanced column sums make the aggregate quantity $M_{1}(t)+M_{2}(t)$ behave like a scalar affine interpretation, even though each coordinate may mix both inputs through off-diagonal coefficients. Strict componentwise decrease forces strict decrease of this aggregate sum, so Theorem 4.2 applies to the aggregate projection. This is the first matrix-style barrier in the artifact whose proof is not a single-coordinate projection theorem, but it still stops short of arbitrary mixed matrix interpretations. $\square$
@@
satisfies the affine equations of Theorem 4.2 with positive wrapper coefficients and an unbounded pump in that weighted projection. Then the duplicating step cannot be strictly oriented componentwise by $M$. In particular, singleton weights recover the tracked-coordinate argument, while the all-ones vector recovers aggregate-sum arguments. Machine-checked in Meta/MatrixBarrierFunctional.lean as no_matrixFunctional_orients_dup_step_of_componentwise_pump together with the successor-pump and wrap-pump corollaries and the KO7 specialization.
Proof sketch. Strict componentwise decrease forces strict decrease of every nonzero natural-weighted scalar projection. Apply this to $\lambda\cdot M$. By hypothesis, that projection already satisfies the scalar affine interface of Theorem 4.2. The contradiction therefore reduces to the scalar affine barrier after one projection step. This unifies the tracked-coordinate proof shape and the balanced aggregate-sum proof shape under one matrix-side principle: once a fixed scalar projection carries the affine interface, strict componentwise orientation is impossible. $\square$
@@
Let $(T,\mathrm{base},\mathrm{succ},\mathrm{wrap},\mathrm{recur})$ be a step-duplicating schema. Let $\mu:T\to A$ be any direct orienter into a codomain $A$, let $R$ be the chosen strict order on $A$, and let $\pi:A\to\mathbb{N}$ be a tracked scalar projection. If every strict $R$-decrease forces strict decrease of $\pi$, and if strict decrease of $\pi\circ\mu$ on the duplicating step is already ruled out by a scalar barrier theorem, then $R$ itself cannot orient the duplicating step uniformly. Machine-checked in Meta/ScalarProjectionBarrier.lean as no_orients_dup_step_of_scalar_projection, with concrete rederivations for tracked componentwise pairs, weighted functional matrices, and balanced mixed-coordinate pairs.
Proof sketch. Suppose $R$ oriented the duplicating step uniformly. By hypothesis, every such $R$-decrease forces strict decrease of the projected scalar $\pi$. Then the projected scalar family $\pi\circ\mu$ would also orient the duplicating step uniformly, contradicting the already available scalar barrier theorem. The value of the theorem is conceptual: it states once the common proof shape shared by the tracked-component, weighted-functional, and balanced mixed-coordinate matrix barriers. $\square$
@@
such that $\succ$ satisfies the standard variable condition: if $x\succ y$, then every variable occurs in $y$ at most as often as in $x$. Then $\succ$ cannot orient the duplicating step from left to right. Machine-checked in Meta/SymbolicComparatorBarrier.lean as not_orients_dup_rule.
Proof sketch. The variable $s$ occurs once on the source side and twice on the target side. Any order satisfying the variable condition must therefore reject this orientation immediately. This widens the negative story beyond numeric measures: direct KBO-style symbolic comparators already fail at the rule schema before any recursive path descent or modular transformation is introduced. $\square$
@@
No KBO-style order satisfying the standard variable condition can orient the KO7 rec_succ rule, and hence no such KBO-style order can orient the KO7 root relation by decreasing every root step. Machine-checked in Meta/KBO_Impossible.lean as no_kbo_orients_ko7_rec_succ.
Proof sketch. Instantiate Theorem 4.14 with the KO7 duplicating rule schema. The corollary adds no new proof mechanism; it simply records the KBO-facing consequence under a theorem name reviewers can cite directly. $\square$
This is a KO7-facing corollary extracted from the symbolic variable-condition barrier, not a formalization of generic Knuth–Bendix order metatheory over arbitrary signatures.
Note on the vector / pair extensions. The fixed-dimension componentwise barrier and the lexicographic dimension-$2$ barrier (Theorems 4.9 and 4.10) extend the scope of applicability of the barrier beyond scalar measures, but the core obstruction mechanism remains the affine tracked-component argument of Theorem 4.2. The balanced mixed-coordinate theorem (Theorem 4.11) goes one step further: it allows off-diagonal coordinate mixing, but only inside a balanced regime where the aggregate sum of the two coordinates still yields a scalar affine projection. The weighted functional theorem (Theorem 4.12) abstracts both moves: any nonzero fixed scalar projection with an affine interface inherits the same barrier. The scalar-projection meta-theorem (Theorem 4.13) then states that proof shape once and for all. So the matrix story now has four layers, tracked-coordinate, balanced mixed-coordinate, weighted scalar projection, and projection-level unification, while arbitrary mixed matrix interpretations remain outside the present barrier stack.

### 08 Apr 2026 v8

@@ 4.2 Affine / linear constructor-local measures
Formalized in Meta/StepDuplicatingSchema.lean.
Both alternatives entail unbounded attainable values: the successor clause pumps directly, while the wrap/base clause pumps through the left-nested recurrence
$x_{k+1}=\alpha_{w}+\beta_{w}x_{k}+\gamma_{w}M(\mathrm{base}),$
that is, through wrapIter with the left argument iterated and the right argument frozen at $\mathrm{base}$. With $\beta_{w}\geq 1$ and positive constant term $\alpha_{w}+\gamma_{w}M(\mathrm{base})\geq 1$, this yields unbounded growth.
Fix $b=n=\mathrm{base}$ and pump $s$. Write
@@
Thus the generic affine contradiction bound is exact for this family. Intuitively, the canonical family makes the duplicated wrapper contribution grow exactly like the pumped payload $k$, while the source still only sees the fixed counter coefficient $\delta_{r}$; the contradiction therefore flips precisely at the cutoff $k=\delta_{r}$. Formalized as affineThresholdMeasure_exact_cutoff.
@@
with $\beta_{w},\gamma_{w}\geq 1$. If $M$ has unbounded attainable values, then $M$ does not strictly orient the duplicating step uniformly. The same conclusion follows from either a positive successor pump or a positive wrap/base pump, exactly as in Theorem 4.2. See Meta/QuadraticBarrier.lean (no_quadratic_counter_orients_dup_step_of_unbounded and the successor- and wrap-pump corollaries).
The argument follows the same freeze-and-pump pattern. Fix $b=n=\mathrm{base}$ and pump the step argument $s$. The rewrite source is
@@
formalized in Meta/QuadraticCrossTermBarrier.lean as CrossTermBoundedAtBase. Then $M$ does not strictly orient the duplicating step uniformly. In particular, the same conclusion follows from either a positive successor pump or a positive wrap/base pump together with this bounded-coupling condition. Proved in Meta/QuadraticCrossTermBarrier.lean (no_cross_quadratic_orients_dup_step_of_unbounded and KO7 specializations).
@@
By contrast, the successful nonlinear witness $W$ lies outside this bounded-coupling fragment: its escape uses broader cross-variable coupling together with successor non-transparency, so the present base-point dominance condition is not the mechanism that certifies it.
Again fix $b=n=\mathrm{base}$ and pump the step argument $s$. The source coefficient in the pump variable becomes $\gamma_{r}+\zeta_{r}\,M(\mathrm{succ}(\mathrm{base}))$, while the target coefficient becomes $\beta_{w}+\gamma_{w}(\gamma_{r}+\zeta_{r}\,M(\mathrm{base}))$. The bounded-coupling hypothesis says the target coefficient still dominates the source coefficient by at least one. Once the pumped step argument exceeds the frozen constant part, the target overtakes the source and strict decrease is impossible. So the barrier reaches one step beyond the pure counter-square class: some step-counter coupling is still excluded, provided the wrapper gain remains dominant at the base point. $\square$
In the rest of this subsection, the bounded cross-term, bounded multilinear, and generalized bounded-polynomial barriers should be read as exclusion theorems for the frozen-base-dominated subclass, not for the entire syntactic family. The dominance predicate is the load-bearing hypothesis in each statement and is summarized again in Table 4.
For the multilinear theorem, if $B_{0}=M(\mathrm{base})$ and $S_{0}=M(\mathrm{succ}(\mathrm{base}))$, the frozen dominance condition is
$\mathrm{sourceCoeff}+1\leq\mathrm{targetCoeff},$
where $\mathrm{sourceCoeff}$ is the total coefficient of the pumped variable $M(s)$ in the frozen source polynomial $M(\mathrm{recur}(\mathrm{base},s,\mathrm{succ}(\mathrm{base})))$, and $\mathrm{targetCoeff}$ is the coefficient of $M(s)$ in the frozen target polynomial $M(\mathrm{wrap}(s,\mathrm{recur}(\mathrm{base},s,\mathrm{base})))$. In Lean this is the explicit predicate MultilinearDominatedAtBase; once $b$ and $n$ are frozen, every multilinear monomial becomes affine in the pumped variable.
@@
where $\mathcal{T}$ is a finite table of multilinear monomials, each variable appearing at most once per monomial. Assume $M$ has unbounded attainable values and that the frozen dominance condition $\mathrm{sourceCoeff}+1\leq\mathrm{targetCoeff}$ holds. Then $M$ does not strictly orient the duplicating step uniformly. The same conclusion follows from either a positive successor pump or a positive wrap/base pump under the same dominance condition. The Lean development is Meta/MultilinearBarrier.lean (no_multilinear_orients_dup_step_of_unbounded, successor-pump, wrap-pump, and KO7 corollaries).
Proof idea. Freeze $b=\mathrm{base}$ and pump the step argument $s$. Each multilinear monomial becomes affine in $M(s)$ once the other arguments are frozen, so the entire finite monomial table collapses to a constant part plus a frozen coefficient times $M(s)$ on both the source and target sides. The explicit dominance condition says the wrapper-side coefficient still exceeds the frozen source coefficient by at least one. Once $M(s)$ exceeds the frozen constant part, the target overtakes the source and strict decrease becomes impossible. A large finite multilinear family therefore still fails as long as its pumped-variable contribution remains frozen and dominated in this sense. $\square$
@@
Let $(T,\mathrm{base},\mathrm{succ},\mathrm{wrap},\mathrm{recur})$ be a step-duplicating schema. Let $M:T\to\mathbb{N}$ satisfy constructor-local affine equations for $\mathrm{succ}$ and $\mathrm{wrap}$, and let the recursor be given by a finite polynomial table in the three tracked scalar arguments $M(b)$, $M(s)$, and $M(n)$, with arbitrary natural exponents allowed in each monomial. Assume $M$ has unbounded attainable values and that, after freezing $b=\mathrm{base}$ and comparing the source counter $\mathrm{succ}(\mathrm{base})$ with the target counter $\mathrm{base}$, the frozen target-side polynomial eventually dominates the frozen source-side polynomial as the pumped step value grows. Then $M$ does not strictly orient the duplicating step uniformly. The same conclusion follows from either a certified successor pump or a certified wrap pump under the same frozen dominance condition. Equivalently, any successful orienter inside this bounded direct-polynomial family must violate the frozen base-dominance condition. Formalized as no_polynomial_orients_dup_step_of_unbounded, no_global_step_orientation_polynomial_of_succ_pump, and polynomial_escape_requires_failure_of_base_dominance in Meta/PolynomialBarrierGeneral.lean.
Explicitly, the dominance condition is
$\exists K\,\forall X\geq K,\;P_{\mathrm{src}}(X)\leq P_{\mathrm{tgt}}(X),$
where $P_{\mathrm{src}}(X)$ is the one-variable polynomial obtained from $M(\mathrm{recur}(\mathrm{base},s,\mathrm{succ}(\mathrm{base})))$ by freezing $M(b)=M(\mathrm{base})$ and replacing $M(s)$ by $X$, and $P_{\mathrm{tgt}}(X)$ is the corresponding frozen target polynomial for $M(\mathrm{wrap}(s,\mathrm{recur}(\mathrm{base},s,\mathrm{base})))$. In Lean this predicate is EventuallyDominatedAtBase.
@@
Freeze the non-pumped arguments at the base point and compare the source counter $\mathrm{succ}(\mathrm{base})$ with the target counter $\mathrm{base}$. Each monomial then becomes a one-variable polynomial in the pumped step value, possibly with repeated occurrences of that variable. The explicit dominance hypothesis says that from some cutoff onward the frozen target-side polynomial overtakes the frozen source-side polynomial. Choose the pumped step value beyond that cutoff. The target is then at least the source, so strict decrease fails. The argument extends beyond the multilinear regime because repeated variables are allowed; however, the resulting one-variable growth must remain dominated at the frozen base point. $\square$
Accordingly, Theorem 4.6 should be read as the artifact’s parametric direct bounded-polynomial barrier for the frozen-base-dominated subclass: any direct bounded-degree polynomial algebra with the standard succ/wrap interface and the same frozen base-dominance condition falls under it, regardless of the specific finite recursor monomial table.
Corollary 4.7 (WPO-style bounded polynomial branch barrier).
Let $\succ$ be any direct order on a step-duplicating schema such that every strict comparison $x\succ y$ forces strict descent in a bounded degree-bounded polynomial algebra of Theorem 4.6. Under the same unbounded or certified pump hypotheses and the same frozen base-dominance condition, $\succ$ cannot orient the duplicating step uniformly. In particular, the KO7 root relation admits no such direct WPO-style polynomial-branch orienter. See Meta/WPO_PolynomialBarrier.lean (no_wpoPolynomialDirect_orients_dup_step_of_unbounded, no_global_step_orientation_wpoPolynomialDirect_of_succ_pump, wpoPolynomialDirect_escape_requires_failure_of_base_dominance).
No new polynomial argument is needed. If strict comparison in the candidate order already implies strict descent of the bounded polynomial algebra, then any successful orientation would induce a successful instance of Theorem 4.6. The polynomial barrier therefore transfers immediately through that soundness implication. The scope is intentionally narrow: the corollary captures the direct bounded polynomial-algebra branch used inside WPO-style tool implementations, not generic recursive path descent or the full WPO metatheory. $\square$
@@
Assume moreover that $\gamma_{w}\geq 1$. If $M$ has unbounded attainable values, then $M$ does not strictly orient the duplicating step uniformly. The same conclusion also follows from either a positive successor drift or a positive left-wrapper drift, but these are separate pump-side hypotheses used only to generate unbounded inputs. The barrier proof itself uses the positive offset on the wrapped recursive branch. The Lean proof is no_max_orients_dup_step_of_unbounded in Meta/MaxBarrier.lean, with successor-pump, wrap-pump, and KO7 corollaries in the same module.
Sketch of the max argument. Pump the step argument until the step branch $\gamma_{r}+M(s)$ dominates both the base branch and the counter branch inside the recursor. At that point the source becomes $\alpha_{r}+\gamma_{r}+M(s)$, while the target wraps the same visible branch again through a max-plus wrapper. Because the wrapper keeps a visible copy of the pumped branch, and because the right wrapper branch carries the positive offset $\gamma_{w}\geq 1$ (absorbing the at-most-one-step successor offset $\alpha_{s}$), the target is at least the source plus one extra unit on that branch, so strict decrease fails. The obstruction therefore persists even when aggregation is by $\max$ rather than by addition. $\square$
@@
Any direct arctic-style interpretation with a distinguished finite primary projection satisfying the max-plus interface of Theorem 4.8 is blocked by the same argument. In particular, the KO7 root relation admits no such arctic primary-projection orienter under the corresponding unbounded or certified pump hypotheses. Proved as no_arctic_primary_orients_dup_step_of_unbounded in Meta/ArcticBarrier.lean; KO7 corollary: no_global_step_orientation_arctic_primary_of_unbounded.
Forget the arctic carrier and read only the distinguished finite primary coordinate. By hypothesis that projection already satisfies the max-plus interface, so Theorem 4.8 applies directly. The corollary is a tool-facing instantiation of the scalar max barrier rather than a separate max-algebra proof. $\square$
@@
Fix any finite dimension $d\geq 1$. Let $M:T\to\mathbb{N}^{d}$ be a vector-valued constructor-local measure on a step-duplicating schema, ordered by strict componentwise decrease. Assume one tracked coordinate of $M$ satisfies the affine equations of Theorem 4.2 with positive wrapper coefficients and an unbounded pump in that tracked coordinate. Then the duplicating step cannot be strictly oriented componentwise by $M$. In particular, the same conclusion follows if the tracked coordinate admits either a positive successor pump or a positive wrap/base pump. Formalized in Meta/MatrixBarrierD.lean (no_matrixD_orients_dup_step_of_componentwise_pump); the earlier dimension-$2$ development is retained in Meta/MatrixBarrier2.lean as the concrete low-dimensional worked instance.
Strict componentwise orientation implies strict decrease in every coordinate, hence in the tracked one. Project the vector-valued measure to that tracked coordinate and apply Theorem 4.2; the contradiction is immediate. The result covers arbitrary fixed dimension under strict componentwise order, but still only through one tracked affine coordinate, not through arbitrary mixed matrix interactions. $\square$
@@
Let $M:T\to\mathbb{N}^{2}$ be a pair-valued constructor-local measure on a step-duplicating schema, ordered lexicographically by the first component. Assume the first component of $M$ satisfies the affine equations of Theorem 4.2 with positive wrapper coefficients and an unbounded pump in that first component. Then the duplicating step cannot be strictly oriented lexicographically by $M$. In particular, the same conclusion follows if the first component admits either a positive successor pump or a positive wrap/base pump. See Meta/MatrixBarrierLex.lean (no_matrix2_lex_orients_dup_step_of_unbounded_primary and KO7 specializations).
Lexicographic decrease implies that the first component of the target is at most the first component of the source. Pump the tracked first component one unit beyond the scalar affine bound. The first component of the duplicating-step target then exceeds the first component of the source, so lexicographic decrease is impossible regardless of what the second component does. $\square$
@@
Let $M:T\to\mathbb{N}^{2}$ be a pair-valued constructor-local measure on a step-duplicating schema, ordered by strict componentwise decrease. Assume each constructor acts by a full mixed $2\times 2$ linear map on the two coordinates, and that the two column sums of each such map agree. Then the aggregate coordinate sum $M_{1}+M_{2}$ is an affine scalar measure. If that aggregate sum has an unbounded pump and the aggregate wrapper coefficients are positive, then the duplicating step cannot be strictly oriented componentwise by $M$. The Lean proof is no_matrixMix2_orients_dup_step_of_sum_pump in Meta/MatrixBarrierMix2.lean.
Aggregate-sum reduction. Balanced column sums make the aggregate quantity $M_{1}(t)+M_{2}(t)$ behave like a scalar affine interpretation, even though each coordinate may mix both inputs through off-diagonal coefficients. Strict componentwise decrease forces strict decrease of this aggregate sum, so Theorem 4.2 applies to the aggregate projection. Unlike the earlier matrix barriers, the proof here is not a single-coordinate projection theorem; it still stops short of arbitrary mixed matrix interpretations however. $\square$
@@
satisfies the affine equations of Theorem 4.2 with positive wrapper coefficients and an unbounded pump in that weighted projection. Then the duplicating step cannot be strictly oriented componentwise by $M$. In particular, singleton weights recover the tracked-coordinate argument, while the all-ones vector recovers aggregate-sum arguments. Proved in Meta/MatrixBarrierFunctional.lean (no_matrixFunctional_orients_dup_step_of_componentwise_pump), with successor-pump and wrap-pump corollaries.
Strict componentwise decrease forces strict decrease of every nonzero natural-weighted scalar projection. Apply this to $\lambda\cdot M$. By hypothesis, that projection already satisfies the scalar affine interface of Theorem 4.2. The contradiction therefore reduces to the scalar affine barrier after one projection step, unifying the tracked-coordinate and balanced aggregate-sum proof shapes under one principle: once a fixed scalar projection carries the affine interface, strict componentwise orientation is impossible. $\square$
@@
Let $(T,\mathrm{base},\mathrm{succ},\mathrm{wrap},\mathrm{recur})$ be a step-duplicating schema. Let $\mu:T\to A$ be any direct orienter into a codomain $A$, let $R$ be the chosen strict order on $A$, and let $\pi:A\to\mathbb{N}$ be a tracked scalar projection. If every strict $R$-decrease forces strict decrease of $\pi$, and if strict decrease of $\pi\circ\mu$ on the duplicating step is already ruled out by a scalar barrier theorem, then $R$ itself cannot orient the duplicating step uniformly. Formalized as no_orients_dup_step_of_scalar_projection in Meta/ScalarProjectionBarrier.lean, with concrete rederivations for tracked componentwise pairs, weighted functional matrices, and balanced mixed-coordinate pairs.
Suppose for contradiction that $R$ oriented the duplicating step uniformly. Every such $R$-decrease forces strict decrease of $\pi$, so the projected scalar family $\pi\circ\mu$ would also orient the duplicating step uniformly, contradicting the already available scalar barrier theorem. The result captures in one statement the common proof shape shared by the tracked-component, weighted-functional, and balanced mixed-coordinate matrix barriers. $\square$
@@
such that $\succ$ satisfies the standard variable condition: if $x\succ y$, then every variable occurs in $y$ at most as often as in $x$. Then $\succ$ cannot orient the duplicating step from left to right. See Meta/SymbolicComparatorBarrier.lean (not_orients_dup_rule).
The variable $s$ occurs once on the source side and twice on the target side. Any order satisfying the variable condition must reject this orientation immediately, widening the negative story beyond numeric measures: direct KBO-style symbolic comparators already fail at the rule schema before any recursive path descent or modular transformation is introduced. $\square$
@@
No KBO-style order on concrete KO7 traces, once restricted along the instantiated duplicating schema and assumed to satisfy the standard variable condition there, can orient the KO7 rec_succ rule. Consequently, no such KBO-style order can orient the KO7 root relation by decreasing every root step. The formalized concrete bridge theorem is no_kbo_orients_ko7_rec_succ_trace in Meta/KBO_Impossible.lean; the underlying abstract schema-level obstruction remains no_kbo_orients_ko7_rec_succ.
Instantiate Theorem 4.15 with the KO7 duplicating rule schema. No new proof mechanism is needed; the corollary records the KBO-facing consequence under a theorem name reviewers can cite directly. $\square$
The corollary is extracted from the symbolic variable-condition barrier and does not formalize generic Knuth–Bendix order metatheory over arbitrary signatures.
Architectural note on Meta/KBO_Impossible.lean. The cited Lean module is intentionally a thin paper-facing renaming layer. The substantive obstruction lives in Meta/SymbolicComparatorBarrier.lean (not_orients_dup_rule) and is reused unchanged. The KBO_Impossible module contributes only a type alias KBOStyleOrder:=VariableConditionOrder, two one-line forwarding theorems under KBO-facing names, and the new concrete-trace bridge no_kbo_orients_ko7_rec_succ_trace that lifts the schema-level statement to literal Trace terms via the instantiate map. Readers searching the artifact for the actual proof of the variable-condition obstruction should look at Meta/SymbolicComparatorBarrier.lean, not at this file.
Note on the vector / pair extensions. The fixed-dimension componentwise barrier and the lexicographic dimension-$2$ barrier (Theorems 4.10 and 4.11) push the barrier beyond scalar measures, but the obstruction is still the affine tracked-component argument of Theorem 4.2. The balanced mixed-coordinate theorem (Theorem 4.12) allows off-diagonal coordinate mixing, but only in a balanced regime where the aggregate sum of the two coordinates still yields a scalar affine projection. The weighted functional theorem (Theorem 4.13) abstracts both moves: any nonzero fixed scalar projection with an affine interface inherits the same barrier. The scalar-projection meta-theorem (Theorem 4.14) states that proof shape once and for all. The matrix story therefore has four layers: tracked-coordinate, balanced mixed-coordinate, weighted scalar projection, and projection-level unification. Arbitrary mixed matrix interpretations remain outside the present barrier stack.